# InsurTech innovation using natural language processing

### v2

@@ 1 Introduction
Despite its clear potential and extensive applications in other fields, such as consumer behavior [101] and banking [142], NLP remains underexplored in actuarial science literature due to limited unstructured textual data availability [110, 108, 147, 76, 145, 144, 129, 148, 86, 143]. This study aims to bridge this gap by providing a comprehensive overview of NLP techniques with a focus on their mathematical foundations. We explore essential methods such as sentiment analysis, topic modeling, language models, and embeddings, and discuss their relevance to various insurance applications. By leveraging real-world InsurTech datasets, we demonstrate how these NLP techniques can:
@@ 2 A Brief Survey of NLP
In the book Foundation of Statistical Natural Language Processing, [111] provide a comprehensive overview of statistical NLP methods developed in the 1990s and before, including the n-gram models, Markov Models, and probability-based lexicon building. Additionally, topic modeling techniques such as Latent Semantic Analysis (LSA) [92] and Probabilistic Latent Semantic Analysis (PLSA) [104] were proposed around the same decade, the 1990s. The well-known Recurrent Neural Network (RNN) [95] also became popular and remains an important model architecture to this day.
In the 2000s, a widely used topic modeling method, Latent Dirichlet Analysis, was proposed by [80]. Researchers also focused on representing words as vectors for semantic and syntactic analysis, leading to the development of word embeddings like word2vec [116, 117] and GloVe [119] between 2012 and 2016. These word embeddings open another level of opportunities to apply natural languages in statistical studies as well as deep learning projects.
In 2017, [138] propose the Attention Mechanism in their paper Attention Is All You Need that reshaped the landscape of NLP from RNN to the Transformer architecture. This breakthrough paved the way for the development of GPT [124], BERT [93], GPT-2 [125], GPT-3 [85], GPT-X (later developments) , and other pre-trained deep learning networks like Llama [136, 96], Gemma [134, 135], Qwen [146, 139, 123] and DeepSeek [90, 91, 99], which have pushed NLP to unprecedented levels of performance and ushered in the era of ChatGPT.
In addition, there are alternative architectures that differ from the Transformer and may have the potential to further advance Large Language Models (LLMs), to list a few, State Space Models (SSMs) like Mamba [98], and newly designed Recurrent Models like RWKV [118].
@@ 3.1.2 Lexicon Normalization: Tokenizing, Stemming, Lemmatization
In the subsequent phase of text preprocessing, various linguistic techniques are applied to the corpus to normalize the texts. Tokenizing involves separating the corpus into words or phrases that are treated as tokens. In some cases, these tokens can be encoded into numerical representations, allowing for the numeric representation of sentences, paragraphs, and even entire documents through further preprocessing. The collection of distinct tokens is referred to as the dictionary of a document collection. To minimize the size of the dictionary, techniques such as Stemming and Lemmatization are applied. Stemming involves obtaining the root form of words. Usually, a word has multiple forms, such as the plural ‘s’ of nouns, the ‘ing’ of verbs, or other affixes, and reducing these forms to a common root simplifies the complexity of the corpus while preserving meaning. The widely used algorithm, the Porter-Stemmer, was introduced in 1980 by [121], which employs a set of production rules for the iterative transformation of English words into their stems. Lemmatization is another method that can be applied to the corpus to reduce complexity by mapping verb forms to the infinitive tense and nouns to the singular form. Lemmatization uses the morphological information of each word to simplify the word into its base form.
@@ 3.2 Embedding
The exploration of word embedding commenced with the Vector Space Model [130] in 1975. In 1988, a method that uses Latent Semantic Analysis was proposed by [92]. Subsequently, in the early 2000s, [78] invented the embedding layer in their Neural Network Language Model paper. Then, word embedding relies on contextual information and word relationships, with influential approaches like word2vec [116, 117], GloVe [119], GPT [124], BERT [93] and its variants. These methods have revolutionized word embedding and found applications beyond NLP, extending to fields like computer vision and recommender systems development.
@@ 3.2.1 Vector Space Model
The Vector Space Model holds significance in NLP despite its age, since it is the first proposal to embed words into vectors that allows the algebraic representation for textual data. Further, the vector representation of terms serves as the cornerstone for incorporating textual data into deep learning models. [130] propose representing a document $d_{j}$, $j=1,\dots,n$, using an m-dimensional vector with selected terms $t_{i}$, $i=1,\dots,m$, within a corpus $D$ containing $n$ documents. The definition of a term, often dictated by the application, can encompass single words, keywords, or longer phrases. The value within the vector is the weight of each term in the document with terms weighted based on their importance. [130] employ the Term Frequency Inverse Document Frequency (TF-IDF) method to weigh the terms within the document. The mathematical formulation of TF-IDF is articulated through a series of equations that define its components and provide an intuitive understanding. The Term Frequency (TF) refers to the relative frequency of a specific term within a document and can be represented as
@@
There are multiple ways to define $TF(t,d)$ and $IDF(t,D)$, along with their combined weighting schemes TFIDF(t,d,D)TF-IDF(t,d,D), each grounded in different intuitions for representing term weight within a document. Additionally, Positive Pointwise Mutual Information (PPMI) offers an alternative scheme of term weighting using term-term matrices, where vector dimensions correspond to terms rather than documents. PPMI draws on the intuition that the strength of association between two terms should be measured by how much more frequently they co-occur in a corpus than would be expected by chance. By leveraging pointwise mutual information, which quantifies this co-occurrence relative to an independence assumption, [89] define alternative intuitions to represent term weight within surrounding terms. These intuitions provide valuable insights that enhance the interpretability of NLP models. This interpretability is particularly critical in the field of insurance, where understanding the rationale behind model predictions is essential for risk assessment, regulatory compliance, and building trust with stakeholders. By ensuring that the term weighting schemes and their combinations are intuitively grounded, insurance practitioners can better understand and explain the model’s decision-making processes.
@@ 3.2.2 Statistical Language Model
As discussed in the previous section, documents can be represented as vectors of terms in the Vector Space Model. However, a significant drawback of this approach is that it can suffer from high-dimensionality, particularly when dealing with a large corpus. To address this issue, [92] propose Latent Semantic Analysis (LSA), which assumes that terms with similar meanings usually appear in similar contexts within documents. Based on this assumption, LSA analyzes relationships between a set of documents and their terms by producing a set of concepts that capture the underlying structure of the documents and terms.
@@
Therefore, $m$ rows, vectors corresponding to a term, in $A_{D}$ map to $k$ columns of $U_{k}$, which represents a lower-dimensional space. Similarly, the $n$ columns, vectors corresponding to a document, in $A_{D}$ map to $k$ columns of $V_{k}$. This lower-dimensional approximation condenses the higher-dimensional space by retaining essential information and filtering out incidental occurrences of terms and documents. In the original matrix $A_{D}$, each entry represents each term presented in each document. In contrast, in the lower-dimensional representation, only terms related to each document are considered in $U_{k}$. However, it is worth pointing out that these new dimensions do not correspond to comprehensible concepts and, therefore, lose interpretability. In addition, the lower-dimensional approximation is derived purely through mathematical matrix operations without incorporating any consideration of natural language structure or linguistic meaning. Inspired by the early SVD-based approaches, several alternative matrix factorization approaches have been developed, including Probabilistic Latent Semantic Indexing (PLSI) [103], Latent Dirichlet Allocation (LDA) [80], and Non-negative Matrix Factorization (NMF) [107], each introducing different probabilistic and structural assumptions to improve representation quality.
SVD and LSA attempt to establish relationships between terms and documents; however, they overlook the relationships between terms themselves, particularly the order in which terms are presented in the document. Another approach, known as a Language Model (LM), involves representing terms probabilistically to predict a term based on the preceding terms. For example, the $\mathcal{N}$-gram language model predicts a term given the $\mathcal{N}-1$ term, thereby assigning probabilities to entire $\mathcal{N}$ sequences. The $\mathcal{N}$-gram LM is developed based on the idea from [114], known as Markov chains, to predict the next word.
@@ 3.2.3 Neural Language Model
Word2vec embeddings are considered static embeddings, meaning they learn a fixed representation for each term in the corpus. Several other static embedding models have been developed to address specific challenges. For example, fastText [^1] FastText is an open-source library that has 157 languages pre-trained embeddings. See https://fasttext.cc/ [82], an extension of word2vec, tackles the issue of unknown terms—those that appear in a corpus but are unseen during training. Additionally, it addresses word sparsity, where infrequent variations of nouns and verbs may have limited occurrences. FastText overcomes these problems by incorporating subword models, representing each term as a combination of itself and a bag of constituent $\mathcal{N}$-grams, with special boundary symbols added. This allows unknown terms to be represented solely by the sum of their constituent $\mathcal{N}$-grams.
Another widely used static embedding model is GloVe [119], which captures global corpus statistics. Unlike word2vec, which relies on local context windows, GloVe constructs word embeddings based on ratios of probabilities derived from the term-term co-occurrence matrix. This approach effectively combines the strengths of count-based models like PPMI, which leverage corpus-wide statistics, and predictive models like word2vec, which capture linear relationships between terms.
[78] is one of the pioneers in introducing neural LM using a feedforward NN. The objective function remains the same as in the $\mathcal{N}$-gram LM, as described in Expression 6. We denote the feedforward NN as $f(\textbf{E},\textbf{W},\textbf{b},\textbf{U})$, where E represents the embeddings for the input terms; precisely, the matrix of the concatenated $\mathcal{N}$ previous terms’ vector representation $v_{t_{i}},i=1,\dots,\mathcal{N}$. These embeddings can be sourced from word2vec, GloVe, or any other pre-trained embedding models. The parameters W and b correspond to the conventional weight matrix and bias vector of the feedforward NN. Additionally, to distinguish the different neural LM architectures, we introduce the weight matrix U, which connects the hidden layers, $h$, to the output layer. In traditional feedforward NN, U is typically a part of W. Given an appropriate loss function, the feedforward NN is trained to predict the next term based on the $\mathcal{N}$ preceding input terms. This training process not only results in a LM, but also produces a new set of optimized embeddings E, which can serve as improved word representations for the training corpus. Figure 1(a) depicts the architecture of the neural LM using a simple feedforward NN.
@@
Language is inherently sequential, unfolding over time in both written and spoken forms. However, feedforward NN lacks a built-in mechanism to capture this temporal dependency. An alternative deep learning architecture that explicitly models sequential dependencies is the Recurrent Neural Network (RNN) [95], along with its more advanced variants, such as Long Short-Term Memory (LSTM) networks [102].
@@
Expanding on the RNN model architecture, encoder-decoder networks—also known as sequence-to-sequence models—are designed to generate contextually appropriate output sequences of arbitrary length given an input sequence. The fundamental idea behind encoder-decoder networks is the use of an encoder network to process the input sequence and create a contextualized representation, often referred to as the context vector. This representation is then passed to a decoder, which generates a task-specific output sequence. The modern neural encoder-decoder architecture is pioneered by [106], who introduced a convolutional neural network (CNN) encoder paired with a RNN decoder. [88], who coined the term encoder-decoder, and [133] show how to use extended RNNs for both encoder and decoder. A significant advancement in encoder-decoder networks comes with the introduction of the attention mechanism. The concept of providing the generative decoder with a weighted combination of input representations—rather than relying solely on a fixed-size context vector—was first explored by [97] in the context of handwriting recognition. The primary function of attention is to alleviate the bottleneck problem by allowing the decoder to access all hidden states of the encoder rather than just the final one. At its core, attention is a weighted sum of context vectors, with additional mechanisms for computing the weights dynamically based on input relevance.
Building upon attention mechanisms, the transformer architecture [138] introduced a paradigm shift by eliminating recurrence entirely. Unlike RNNs, which process sequences step-by-step, transformers leverage attention mechanisms to model dependencies between terms in parallel, significantly improving efficiency and scalability. A key advantage of transformers is their ability to generate contextual embeddings, where the representation of a term is dynamically adjusted based on its surrounding context—unlike static embeddings such as word2vec, which assign a fixed vector to each word regardless of its usage. The input representation in transformers is more complex compared to RNNs, as it involves separately computing two embeddings: (1) an input token embedding, which represents the meaning of the term, and (2) an input positional embedding, which encodes the term’s position within the sequence. This additional positional information is crucial since transformers, unlike RNNs, do not inherently process input in sequential order. The combination of these embeddings, along with multi-head attention and feedforward layers, allows transformers to capture nuanced contextual relationships, making them the backbone of modern NLP models such as BERT [93] and GPT [85]. At the time of writing this paper, the most recent advancements in NLP largely remain based on encoder-decoder architectures and their variants.

### v3

@@ 1 Introduction
Despite its clear potential and extensive applications in other fields, NLP remains underexplored in actuarial science literature due to limited unstructured textual data availability [Liao2020, Lee2020, zappa_text_2021, baillargeon2021mining, xu_bert-based_2022, xu2023framework, sabban2022automatic, Zhang11112024, cao2024assessing]. This study aims to bridge this gap by providing a comprehensive overview of NLP techniques with a focus on their mathematical foundations. We explore essential methods such as sentiment analysis, topic modeling, language models, and embeddings, and discuss their relevance to various insurance applications. By leveraging real-world InsurTech datasets, we demonstrate how these NLP techniques can:
@@ 2 A Brief Survey of NLP
In the book Foundation of Statistical Natural Language Processing, [ManningBook] provide a comprehensive overview of statistical NLP methods developed in the 1990s and before, including the n-gram models, Markov Models, and probability-based lexicon building. Additionally, topic modeling techniques such as Latent Semantic Analysis (LSA) [Deerwester1990] and Probabilistic Latent Semantic Analysis (PLSA) [Hofmann2001] were proposed around the same decade, the 1990s. The well-known Recurrent Neural Network (RNN) [elman1990finding] also became popular and remains an important model architecture to this day.
In the 2000s, a widely used topic modeling method, Latent Dirichlet Analysis, was proposed by [Blei2003]. Researchers also focused on representing words as vectors for semantic and syntactic analysis, leading to the development of word embeddings like word2vec [Mikolov2013, Mikolov2013a] and GloVe [Pennington2014GloveGV] between 2012 and 2016. These word embeddings open another level of opportunities to apply natural languages in statistical studies as well as deep learning projects.
In 2017, [Vaswani2017] propose the Attention Mechanism in their paper Attention Is All You Need that reshaped the landscape of NLP from RNN to the Transformer architecture. This breakthrough paved the way for the development of GPT [radford2018improving], BERT [devlin2019], GPT-2 [radford2019language], GPT-3 [brown2020language], GPT-X (later developments) , and other pre-trained deep learning networks like Llama [touvronLlamaOpenFoundation2023, grattafioriLlama3Herd2024], Gemma [teamGemmaOpenModels2024, teamGemma2Improving2024], Qwen [yangQwen2TechnicalReport2024, wangQwen2VLEnhancingVisionLanguage2024, qwenQwen25TechnicalReport2025] and DeepSeek [deepseek-aiDeepSeekV2StrongEconomical2024, deepseek-aiDeepSeekV3TechnicalReport2024, guoDeepSeekR1IncentivizesReasoning2025], which have pushed NLP to unprecedented levels of performance and ushered in the era of ChatGPT.
In addition, there are alternative architectures that differ from the Transformer and may have the potential to further advance Large Language Models (LLMs), to list a few, State Space Models (SSMs) like Mamba [guMambaLinearTimeSequence2024], and newly designed Recurrent Models like RWKV [pengRWKVReinventingRNNs2023].
@@ 3.1.2 Lexicon Normalization: Tokenizing, Stemming, Lemmatization
In the subsequent phase of text preprocessing, various linguistic techniques are applied to the corpus to normalize the texts. Tokenizing involves separating the corpus into words or phrases that are treated as tokens. In some cases, these tokens can be encoded into numerical representations, allowing for the numeric representation of sentences, paragraphs, and even entire documents through further preprocessing. The collection of distinct tokens is referred to as the dictionary of a document collection. To minimize the size of the dictionary, techniques such as Stemming and Lemmatization are applied. Stemming involves obtaining the root form of words. Usually, a word has multiple forms, such as the plural ‘s’ of nouns, the ‘ing’ of verbs, or other affixes, and reducing these forms to a common root simplifies the complexity of the corpus while preserving meaning. The widely used algorithm, the Porter-Stemmer, was introduced in 1980 by [Porter1980], which employs a set of production rules for the iterative transformation of English words into their stems. Lemmatization is another method that can be applied to the corpus to reduce complexity by mapping verb forms to the infinitive tense and nouns to the singular form. Lemmatization uses the morphological information of each word to simplify the word into its base form.
@@ 3.2 Embedding
The exploration of word embedding commenced with the Vector Space Model [Salton1975] in 1975. In 1988, a method that uses Latent Semantic Analysis was proposed by [Deerwester1990]. Subsequently, in the early 2000s, [Bengio2001] invented the embedding layer in their Neural Network Language Model paper. Then, word embedding relies on contextual information and word relationships, with influential approaches like word2vec [Mikolov2013, Mikolov2013a], GloVe [Pennington2014GloveGV], GPT [radford2018improving], BERT [devlin2019] and its variants. These methods have revolutionized word embedding and found applications beyond NLP, extending to fields like computer vision and recommender systems development.
@@ 3.2.1 Vector Space Model
The Vector Space Model holds significance in NLP despite its age, since it is the first proposal to embed words into vectors that allows the algebraic representation for textual data. Further, the vector representation of terms serves as the cornerstone for incorporating textual data into deep learning models. [Salton1975] propose representing a document $d_{j}$, $j=1,\dots,n$, using an m-dimensional vector with selected terms $t_{i}$, $i=1,\dots,m$, within a corpus $D$ containing $n$ documents. The definition of a term, often dictated by the application, can encompass single words, keywords, or longer phrases. The value within the vector is the weight of each term in the document with terms weighted based on their importance. [Salton1975] employ the Term Frequency Inverse Document Frequency (TF-IDF) method to weigh the terms within the document. The mathematical formulation of TF-IDF is articulated through a series of equations that define its components and provide an intuitive understanding. The Term Frequency (TF) refers to the relative frequency of a specific term within a document and can be represented as
@@
There are multiple ways to define $TF(t,d)$ and $IDF(t,D)$, along with their combined weighting schemes TFIDF(t,d,D)TF-IDF(t,d,D), each grounded in different intuitions for representing term weight within a document. Additionally, Positive Pointwise Mutual Information (PPMI) offers an alternative scheme of term weighting using term-term matrices, where vector dimensions correspond to terms rather than documents. PPMI draws on the intuition that the strength of association between two terms should be measured by how much more frequently they co-occur in a corpus than would be expected by chance. By leveraging pointwise mutual information, which quantifies this co-occurrence relative to an independence assumption, [church1990word] define alternative intuitions to represent term weight within surrounding terms. These intuitions provide valuable insights that enhance the interpretability of NLP models. This interpretability is particularly critical in the field of insurance, where understanding the rationale behind model predictions is essential for risk assessment, regulatory compliance, and building trust with stakeholders. By ensuring that the term weighting schemes and their combinations are intuitively grounded, insurance practitioners can better understand and explain the model’s decision-making processes.
@@ 3.2.2 Statistical Language Model
As discussed in the previous section, documents can be represented as vectors of terms in the Vector Space Model. However, a significant drawback of this approach is that it can suffer from high-dimensionality, particularly when dealing with a large corpus. To address this issue, [Deerwester1990] propose Latent Semantic Analysis (LSA), which assumes that terms with similar meanings usually appear in similar contexts within documents. Based on this assumption, LSA analyzes relationships between a set of documents and their terms by producing a set of concepts that capture the underlying structure of the documents and terms.
@@
Therefore, $m$ rows, vectors corresponding to a term, in $A_{D}$ map to $k$ columns of $U_{k}$, which represents a lower-dimensional space. Similarly, the $n$ columns, vectors corresponding to a document, in $A_{D}$ map to $k$ columns of $V_{k}$. This lower-dimensional approximation condenses the higher-dimensional space by retaining essential information and filtering out incidental occurrences of terms and documents. In the original matrix $A_{D}$, each entry represents each term presented in each document. In contrast, in the lower-dimensional representation, only terms related to each document are considered in $U_{k}$. However, it is worth pointing out that these new dimensions do not correspond to comprehensible concepts and, therefore, lose interpretability. In addition, the lower-dimensional approximation is derived purely through mathematical matrix operations without incorporating any consideration of natural language structure or linguistic meaning. Inspired by the early SVD-based approaches, several alternative matrix factorization approaches have been developed, including Probabilistic Latent Semantic Indexing (PLSI) [10.1145/312624.312649], Latent Dirichlet Allocation (LDA) [Blei2003], and Non-negative Matrix Factorization (NMF) [lee1999learning], each introducing different probabilistic and structural assumptions to improve representation quality.
SVD and LSA attempt to establish relationships between terms and documents; however, they overlook the relationships between terms themselves, particularly the order in which terms are presented in the document. Another approach, known as a Language Model (LM), involves representing terms probabilistically to predict a term based on the preceding terms. For example, the $\mathcal{N}$-gram language model predicts a term given the $\mathcal{N}-1$ term, thereby assigning probabilities to entire $\mathcal{N}$ sequences. The $\mathcal{N}$-gram LM is developed based on the idea from [markov1913essai], known as Markov chains, to predict the next word.
@@ 3.2.3 Neural Language Model
Word2vec embeddings are considered static embeddings, meaning they learn a fixed representation for each term in the corpus. Several other static embedding models have been developed to address specific challenges. For example, fastText [^1] FastText is an open-source library that has 157 languages pre-trained embeddings. See https://fasttext.cc/ [bojanowski-etal-2017-enriching], an extension of word2vec, tackles the issue of unknown terms—those that appear in a corpus but are unseen during training. Additionally, it addresses word sparsity, where infrequent variations of nouns and verbs may have limited occurrences. FastText overcomes these problems by incorporating subword models, representing each term as a combination of itself and a bag of constituent $\mathcal{N}$-grams, with special boundary symbols added. This allows unknown terms to be represented solely by the sum of their constituent $\mathcal{N}$-grams.
Another widely used static embedding model is GloVe [Pennington2014GloveGV], which captures global corpus statistics. Unlike word2vec, which relies on local context windows, GloVe constructs word embeddings based on ratios of probabilities derived from the term-term co-occurrence matrix. This approach effectively combines the strengths of count-based models like PPMI, which leverage corpus-wide statistics, and predictive models like word2vec, which capture linear relationships between terms.
[Bengio2001] is one of the pioneers in introducing neural LM using a feedforward NN. The objective function remains the same as in the $\mathcal{N}$-gram LM, as described in Expression 6. We denote the feedforward NN as $f(\textbf{E},\textbf{W},\textbf{b},\textbf{U})$, where E represents the embeddings for the input terms; precisely, the matrix of the concatenated $\mathcal{N}$ previous terms’ vector representation $v_{t_{i}},i=1,\dots,\mathcal{N}$. These embeddings can be sourced from word2vec, GloVe, or any other pre-trained embedding models. The parameters W and b correspond to the conventional weight matrix and bias vector of the feedforward NN. Additionally, to distinguish the different neural LM architectures, we introduce the weight matrix U, which connects the hidden layers, $h$, to the output layer. In traditional feedforward NN, U is typically a part of W. Given an appropriate loss function, the feedforward NN is trained to predict the next term based on the $\mathcal{N}$ preceding input terms. This training process not only results in a LM, but also produces a new set of optimized embeddings E, which can serve as improved word representations for the training corpus. Figure 1(a) depicts the architecture of the neural LM using a simple feedforward NN.
@@
Language is inherently sequential, unfolding over time in both written and spoken forms. However, feedforward NN lacks a built-in mechanism to capture this temporal dependency. An alternative deep learning architecture that explicitly models sequential dependencies is the Recurrent Neural Network (RNN) [elman1990finding], along with its more advanced variants, such as Long Short-Term Memory (LSTM) networks [10.1162/neco.1997.9.8.1735].
@@
Expanding on the RNN model architecture, encoder-decoder networks—also known as sequence-to-sequence models—are designed to generate contextually appropriate output sequences of arbitrary length given an input sequence. The fundamental idea behind encoder-decoder networks is the use of an encoder network to process the input sequence and create a contextualized representation, often referred to as the context vector. This representation is then passed to a decoder, which generates a task-specific output sequence. The modern neural encoder-decoder architecture is pioneered by [kalchbrenner-blunsom-2013-recurrent], who introduced a convolutional neural network (CNN) encoder paired with a RNN decoder. [cho-etal-2014-learning], who coined the term encoder-decoder, and [NIPS2014_a14ac55a] show how to use extended RNNs for both encoder and decoder. A significant advancement in encoder-decoder networks comes with the introduction of the attention mechanism. The concept of providing the generative decoder with a weighted combination of input representations—rather than relying solely on a fixed-size context vector—was first explored by [graves2013generating] in the context of handwriting recognition. The primary function of attention is to alleviate the bottleneck problem by allowing the decoder to access all hidden states of the encoder rather than just the final one. At its core, attention is a weighted sum of context vectors, with additional mechanisms for computing the weights dynamically based on input relevance.
Building upon attention mechanisms, the transformer architecture [Vaswani2017] introduced a paradigm shift by eliminating recurrence entirely. Unlike RNNs, which process sequences step-by-step, transformers leverage attention mechanisms to model dependencies between terms in parallel, significantly improving efficiency and scalability. A key advantage of transformers is their ability to generate contextual embeddings, where the representation of a term is dynamically adjusted based on its surrounding context—unlike static embeddings such as word2vec, which assign a fixed vector to each word regardless of its usage. The input representation in transformers is more complex compared to RNNs, as it involves separately computing two embeddings: (1) an input token embedding, which represents the meaning of the term, and (2) an input positional embedding, which encodes the term’s position within the sequence. This additional positional information is crucial since transformers, unlike RNNs, do not inherently process input in sequential order. The combination of these embeddings, along with multi-head attention and feedforward layers, allows transformers to capture nuanced contextual relationships, making them the backbone of modern NLP models such as BERT [devlin2019] and GPT [brown2020language]. At the time of writing this paper, the most recent advancements in NLP largely remain based on encoder-decoder architectures and their variants.